OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

My function is f(x) 4x +6 The inverse of it is f-1(x) = x - 6 /4

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

h(x) can be anything.. So, say let us take as h(x) = 2x + 3

OpenStudy (anonymous):

Can you find f(h(x))??

OpenStudy (anonymous):

oops sorry was making a sandwich

OpenStudy (anonymous):

so could you use 3 for x?

OpenStudy (anonymous):

f(3)2x + 3 f(3)6 + 3 f(3)9 What now..?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Sorry, I gotta go now... :) @campbell_st when you are free, please look at this question.. :)

OpenStudy (campbell_st):

well looking at the 2 functions the domain of f(x) will become the range of the inverse f^(-1) and the range of f(x) will become the domain of the inverse f^(-1) that seems to be the idea here... so if you put a value from the domain of f(x) say 1 then f(x) = 10 then substituting that into the inverse will give a solution of 1. hope it makes sense

OpenStudy (anonymous):

so it that supposed to be relating f(h(x)) and h(f(x)) together? also, i dont really get what i'm using the function from #2 for? proving the result will be the same?

OpenStudy (campbell_st):

ok... so do it this way f(x) = 2x h(x) = x + 2 h(f(x) = (2x) + 2 = 2x + 2 now the other way f(h(x)) = 2(x + 2) = 2x + 4 are they the same things.... h(f(x) and f(h(x))...?

OpenStudy (campbell_st):

so in the 1st case h(f(x)) replace x in h(x) with 2x from f(x) then reverse it... do you get the same equation...

OpenStudy (anonymous):

well you ended up with 2x + 2 and 2x +4 so no

OpenStudy (anonymous):

so that's it i guess?